OpenStudy (anonymous):

expand and simplify 9a+3(8-2a)

OpenStudy (anonymous):

9a+24-6a=3a+24

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

9a+3(8-2a) = 9a + 24 - 6a = 3a + 24